If y is not zero and x, y belongs to Q then prove that x/y belongs to Q


1
Expert's answer
2022-01-19T15:28:56-0500


Since "y\in\mathbb{Q},y≠0" and "x\in\mathbb{Q}"

Note that "y≠0,y\in\mathbb{Q}" "\implies\frac{1}{y}\in\mathbb{Q}"

Since "x\in\mathbb{Q}" and "y\in\mathbb{Q}"

"\implies xy\in\mathbb{Q}"

Then "\frac{1}{y}.(xy)=(\frac{1}{y}.x)y=(\frac{x}{y})y=x\in\mathbb{Q}"

Since set of rational numbers is closed under multiplication

"\implies\frac{x}{y}\in\mathbb{Q}"